OpenStudy (anonymous):

841cg = how many dkg

OpenStudy (cuanchi):

\[841 cg *\frac{ 1g }{ 100cg }*\frac{ 1dkg }{ 10g }=\frac{ 841 }{ 1000 } dkg=0.841dkg\]

OpenStudy (supernova_sonntag):

That or you could just move the decimal three places to the left. :)